Suppose the company paid dividend of $40.83. That is, D0 is $40.83. What is the amount of dividend in Year 12? That is, what is D12? Assume that the dividends are expected to grow by 1% each year.

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

In this question we need to determine D12 i.e. dividend to be paid in year 12.

Following information are given:-

D0 = 40.83 g = 1%

As we know that

D1 = D0(1+g)

D1 = 40.83(1+0.01)

D1 = 41.24

Just like above

D2 = D1(1+g)^1

D3 = D1(1+g)^2

.

.

.

.

D12 = D1(1+g)^11

= 41.24(1+.01)^11

= 41.24(1.12)

= 46.19

D12 = 46.19

Note : Figures are rounded off to two decimal points, your answer will be different if you will not consider round off.